Find the pH of a 0.008 M solution of an aluminum compound which dissolves completely to give the hydrated aluminum ion Al(H2O)3+6.

Al(H2O)3+6(aq)+H2O(l)⇌H3O+(aq)+Al(H2O)5(OH)2+(aq),    Ka=1.4×10−5

  • Report your answer with three significant figures.

    Find the pH of a 0.008 M solution of an aluminum compound which dissolves completely to give the hydrated aluminum ion Al(H2O)3+6.

    Al(H2O)3+6(aq)+H2O(l)⇌H3O+(aq)+Al(H2O)5(OH)2+(aq),    Ka=1.4×10−5

  • Report your answer with three significant figures.
Answer #1

Given:-

molar concentration of [Al(H2O)6]3+ = 0.008 M

equilibrium constant of [Al(H2O)6]3+ = 1.4 \times 10-5

As we know that 0.008 M solution of aluminum compound dissolves completely to give the hydrated aluminum ion i.e [Al(H2O)6]3+ then molar concentration of [Al(H2O)6]3+ would be 0.008 M and then equilibrium established between [Al(H2O)6]3+ and solution of aluminum compound i.e[Al(H2O)5 (OH)]2+ which is given as follows:-

[Al(H2O)6]3+ + H2O(l)\rightleftharpoons [Al(H2O)5 (OH)]2+(aq) + H3O+(aq)

0.008 M 0 0

(0.008 - X) X X

As we know that

equilibrium constant of [Al(H2O)6]3+ (Ka) = [ [Al(H2O)5 (OH)]2+ ][H3O+] / [ [Al(H2O)6]3+ ][H2O]

equilibrium constant of [Al(H2O)6]3+ (Ka) =X \timesX / (0.008 - X) \times 1 ( since [H2O] = 1)

neglecting the X in denominator in above expression then we get

equilibrium constant of [Al(H2O)6]3+ (Ka) =  X2 / 0.008

1.4 \times 10-5 =  X2 / 0.008

X2 = 0.008 \times  1.4 \times 10-5

X2 = 0.0112\times 10-5

X2 = 11.2\times 10-8

X = \sqrt{}11.2\times 10-8

X   = 3.3466 \times 10-4

therefore

molar concentration of [H3O+] = 3.3466 \times 10-4

molar concentration of [ [Al(H2O)5 (OH)]2+ ]= 3.3466 \times 10-4

As we know that

pH = - log[H3O+]

pH = - log[3.3466 \times 10-4 ]

pH = - [log(3.3466) + ( -4log(10) ]

pH = - [(0.5246) - ( 4) ]

pH = 4.0000 - 0.5246

pH = 3.4754 (i.e the answer)